# Optimal and Fair Encouragement Policy Evaluation and Learning

### v4

@@ 6 Additional Fairness Constraints and Policy Optimization
4: Augment additional constraints with $\epsilon_{n}$ policy value and constraint slices relative to $\hat{\pi}_{1}$; define an augmented system (where subscripting by $\hat{\mathcal{I}}_{1}$ subindexes the corresponding matrix or vector):
$\tilde{h}_{j^{\prime}}(Q)$ $=\mathbb{E}_{n_{1}}[\{g_{j^{\prime}}(O;\hat{Q}_{1})-g_{j^{\prime}}(O;Q)\}\mid\mathcal{E}_{j}],\;\;\forall j^{\prime}\in\hat{\mathcal{I}}_{1},\;$
$\tilde{h}^{v}(Q)$ $=\mathbb{E}_{n_{1}}[v_{DR}(O;\hat{Q}_{1},\eta_{1})-v_{DR}(O;Q,\eta_{1})]$
$\tilde{M}$ $=[M;M_{\hat{\mathcal{I}}_{1}},\vec{1}],\tilde{h}=[h,\tilde{h},\tilde{h}^{v}]^{\top},\tilde{d}=[\hat{d},\epsilon_{n}\vec{1},\epsilon_{n}]^{\top},\tilde{\mathcal{E}}=[\mathcal{E},\mathcal{E}_{\hat{\mathcal{I}}_{1}},\emptyset]^{\top}$
5: Solve $\min_{Q\in\Delta(\Pi)}\{\hat{V}(Q)\colon\;\;\tilde{M}\tilde{h}(Q)\leq\tilde{d}\}$. Obtain $\hat{Q}^{*}_{2}$ by running Algorithm˜1 on $\mathcal{D}_{2}$ with $\text{REDFAIR}(\mathcal{D},\tilde{g},\tilde{\mathcal{E}},\tilde{M},\tilde{d},\eta_{2})$.
Optimization. Ultimately, the optimization is solved with sampled and estimated moments. Define the integrand of the constrained, weighted empirical risk minimization as $v_{(\cdot)}(O;\pi_{\beta},\eta)=\left|\psi_{(\cdot)}(O;\eta)\right|\ell(f_{\beta}(X),\operatorname{sgn}(\psi_{(\cdot)}(O;\eta))).$ Our estimate of the objective function is therefore
$V_{(\cdot)}(Q)=\mathbb{E}[\left|\psi_{(\cdot)}\right|\ell(f_{\beta},\operatorname{sgn}(\psi_{(\cdot)}))]=\mathbb{E}_{\pi_{\beta}\sim Q}[v_{(\cdot)}(O;\pi_{\beta},\eta)].$
Note that for the rest of our discussions of algorithms for constrained policy optimization, we overload notation and use $V_{(\cdot)}(Q)$ to refer to policy regret, as above. The optimal policies are the same for regret and for value. We obtain the sample estimator $\hat{V}_{(\cdot)}(Q)$ and sample constraint moments $\hat{h}(Q)$ analogously. We also add a feasibility margin $\epsilon_{k}$ that depends on concentration of the estimated constraints, so the sampled constraint vector is $\hat{d}_{k}=d_{k}+\epsilon_{k},$ for all $k$. We seek an approximate saddle point so that the constrained solution is equivalent to the Lagrangian,
$\hat{L}(Q,{\lambda})=\hat{V}(Q)+{\lambda}^{\top}({M}\hat{{h}}(Q)-\hat{{d}}),\qquad\min_{Q\in\Delta(\Pi)}\{\hat{V}(Q)\colon M\hat{h}(Q)\leq\hat{d}\}=\min_{Q\in\Delta(\Pi)}\max_{{\lambda}\in\mathbb{R}_{+}^{K}}\hat{L}(Q,\lambda).$
$\textstyle\underset{Q\in\Delta}{\min}\;\;\underset{{{\lambda}\in\mathbb{R}_{+}^{|\mathcal{X}|},\|{\lambda}\|_{1}\leq B}}{\max}\hat{L}(Q,{\lambda}),\qquad\underset{{\lambda}\in\mathbb{R}_{+}^{|\mathcal{X}|},\|{\lambda}\|_{1}\leq B}{\max}\;\;\underset{Q\in\Delta}{\min}\;\hat{L}(Q,{\lambda})$
(Agarwal et al., 2018, Theorem 3) gives generalization guarantees on the policy value and constraint violation achieved by the approximate saddle point output by the algorithm. The analysis is generic under rate assumptions on uniform convergence of policy and constraint values. Such a rate $\alpha$ follows from standard analyses in causal inference and is used to set the constraint violation feasibility margin $\epsilon_{k}=O(n^{-\alpha})$.
Assumption 7 (Rate assumption on policy and constraint values).
There exist $C,C^{\prime}\geq 0$ and $\alpha\leq 1/2$ such that $\sup_{Q\in\Delta(\Pi)}\{V(Q;\eta)-\hat{V}(Q;\hat{\eta})\}\leq Cn^{-\alpha}$ and $\varepsilon_{k}=C^{\prime}\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|n_{j}^{-\alpha}$, where $n_{j}$ is the number of data points that fall in $\mathcal{E}_{j}$.
Next, we summarize the optimization algorithm. We play a no-regret (second-order multiplicative weights; Cesa-Bianchi et al., 2007; Steinhardt and Liang, 2014) algorithm (a slight variant of the hedge/exponentiated gradient algorithm; Freund and Schapire, 1997) for the $\lambda$-player while using best-response oracles for the $Q$-player. Full details are in Algorithm˜1. Given $\lambda_{t},$ $\operatorname{BEST}_{\beta}\left({\lambda}_{t}\right)$ computes a best response over $Q$; since the worst-case distribution will place all its weight on one classifier, this step can be implemented by a reduction to cost-sensitive/weighted classification (Beygelzimer and Langford, 2009; Zhao et al., 2012), which we describe in further detail below. Computing the best response over $\operatorname{BEST}_{{\lambda}}(\hat{Q}_{t}))$ selects the most violated constraint. Further details are in Section˜C.2.
Two-stage variance-constrained algorithm. We seek to improve upon this procedure so that we may obtain regret bounds on policy value and fairness constraint violation that exhibit more favorable dependence on the maximal variance over small-variance slices near the optimal policy, rather than worst-case constants over all policies (Chernozhukov et al., 2019; Athey and Wager, 2021). Moreover, using the estimated variance to set constraint feasibility slacks can achieve tighter fairness control.
These challenges motivate the two-stage procedure, described formally in Algorithm˜2 and verbally here. We adapt an out-of-sample regularization scheme developed in (Chernozhukov et al., 2019), which recovers variance-sensitive regret bounds via a small modification to an ERM procedure (and, by extension, policy learning). We split the data into two subsets, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, and first learn nuisance estimators $\hat{\eta}_{1}$ from $\mathcal{D}_{1}$ (possibly with further sample splitting) for use in our policy value and constraint estimates. We run Algorithm˜1 ($\operatorname{REDFAIR}(\mathcal{D}_{1},h,{\mathcal{E}},M,d;\hat{\eta}_{1})$) on data from $\mathcal{D}_{1}$ to estimate the optimal policy distribution $\hat{Q}_{1}$ and the constraint variances at $\hat{Q}_{1}$. We identify the first-stage binding constraints via the index set $\hat{I}_{1}$. Next, we augment the constraint matrix with additional constraints that require feasible policies for the second-stage policy distribution to achieve $\epsilon_{n}$ close policy value and constraint moment values relative to $\hat{Q}_{1}$. Since errors concentrate quickly, this can be viewed as variance regularization. In addition, we set the constraint slacks $\hat{d}\leftarrow d+2\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|\hat{\sigma}^{2}_{j}n^{-\alpha}$ in the second stage using the estimated variance constants from $\hat{Q}_{1}$. This results in tighter control of the fairness constraints. The second stage solves for an approximate saddle point of the augmented system, with objective function and constraints evaluated on $\mathcal{D}_{2}$ and returns $\hat{Q}_{2}.$
Next, we provide a generalization bound on the out-of-sample performance of the policy returned by the two-stage procedure. Importantly, because of our two-stage procedure, the regret of the policy depends on the worst-case variance of near-optimal policies (rather than all policies). Define the function classes $\mathcal{F}_{\Pi}=\{v_{DR}(\cdot,\pi;\eta)\colon\pi\in\Pi,\eta\in\mathcal{F}_{\eta}\}$, $\mathcal{F}_{j}=\{g_{j}(\cdot,\pi;\eta)\colon\pi\in\Pi,\eta\in\mathcal{F}_{\eta}\}$ and the empirical entropy integral $\kappa(r,\mathcal{F})={\inf_{\alpha\geq 0}\{4\alpha+10\int_{\alpha}^{r}\sqrt{\frac{\mathcal{H}_{2}(\epsilon,\mathcal{F},n)}{n}}d\epsilon\}}$, where $H_{2}(\epsilon,\mathcal{F},n)$ is the $L_{2}$ empirical entropy, i.e., log of the $\lVert\cdot\rVert_{2}$ $\epsilon$-covering number. We make a mild assumption of a learnable function class (bounded entropy integral) (Van Der Vaart et al., 1996), which is satisfied by many standard function classes such as linear models, polynomials, kernel regression, and neural networks (Wainwright, 2019).
Assumption 8.
The function classes $\mathcal{F}_{\Pi},\{\mathcal{F}_{j}\}_{j\in\mathcal{J}}$ satisfy that, for any constant $r,\kappa(r,\mathcal{F})\rightarrow 0$ as $n\rightarrow\infty$. The function classes $\{\mathcal{F}_{j}\}_{j\in\mathcal{J}}$ comprise $L_{j}$-Lipschitz contractions of $\pi.$
We assume that we are using doubly robust/orthogonalized estimation as in proposition˜5 and, hence, state our results depending on the estimation error of nuisance vector $\eta$. The next theorem summarizes the out-of-sample performance of the two-stage algorithm of Algorithm˜2, $\hat{Q}_{2}$.
Theorem 4 (Variance-based oracle policy regret).
Suppose that the mean-squared error of the nuisance estimates is upper bounded w.p. $1-\delta/2$ by $\chi_{n,\delta}^{2}$ over the randomness of the nuisance sample: $\max_{l}\{\mathbb{E}[(\hat{\eta}_{l}-\eta_{l})^{2}]\}_{l\in[L]}\coloneqq\chi_{n}^{2}$
Let $v_{DR}^{0}(O;Q)$ denote evaluation with true nuisance functions $\eta_{0}$; define $r=\sup_{Q\in\mathcal{Q}}\sqrt{\mathbb{E}\left[v_{DR}^{0}(O;Q)^{2}\right]}$ and $\epsilon_{n}=\Theta\left(\kappa\left(r,\mathcal{F}_{\Pi}\right)+r\sqrt{\frac{\log(1/\delta)}{n}}\right)$. Moreover, denote an $\epsilon$-regret slice of the policy space:
@@
$\mathcal{Q}_{*}(\epsilon)=\left\{Q\in\Delta[\Pi]:V(Q_{*}^{0})-V(Q)\leq\epsilon,\;\;h(Q_{*}^{0})-h(Q)\leq d+\epsilon\right\}$
Let $\tilde{\epsilon}_{n}=O(\epsilon_{n}+\chi_{n,\delta}^{2})$, and denote the variance of the difference between any two policies in an $\epsilon_{n}$-regret slice, evaluated at the true nuisance quantities:
@@
$\bar{\sigma}_{\mathcal{D}_{2}}^{2}=\sup\;\{\operatorname{Var}\left(v_{DR}^{0}(O;Q)-v_{DR}^{0}\left(O;Q^{\prime}\right)\right)\colon{Q,Q^{\prime}\in\mathcal{Q}_{*}(\tilde{\epsilon}_{n})}\}$
(Define $\bar{\sigma}_{k,\mathcal{D}_{2}}^{2}$ analogously for the variance of constraint moments.) Then, letting ${{\gamma}}(Q):={M}{{h}}(Q)$ denote the constraint values, the policy distribution $Q_{2}$ returned by the out-of-sample regularized ERM satisfies w.p. $1-\delta$ over the randomness of $S$:
@@
$V(\hat{Q}_{2})-V(Q^{*})$ $=\textstyle O(\kappa(\bar{\sigma}_{\mathcal{D}_{2}},\operatorname{conv}(\mathcal{F}_{\Pi}))+\bar{\sigma}_{\mathcal{D}_{2}}{n}^{-\frac{1}{2}}\sqrt{{\log(3/\delta)}}+\chi_{n,\delta}^{2})$
@@
$(\gamma_{k}(\hat{Q}_{2})-d_{k})-(\gamma_{k}(Q^{*})-d_{k})$ $=\textstyle O(\kappa(\bar{\sigma}_{k,\mathcal{D}_{2}},\operatorname{conv}(\mathcal{F}_{j}))+\bar{\sigma}_{k,\mathcal{D}_{2}}{n}^{-\frac{1}{2}}\sqrt{{\log(3/\delta)}}+\chi_{n,\delta}^{2})$
@@
The specific benefits of the two-stage approach are that 1) the constants are improved from being absolute, structure-agnostic bounds to depending on the variance of low-regret policies, which also reflects the improved variance from the use of doubly robust estimation as in proposition˜5, and 2) it allows less conservative satisfaction of the fairness constraint out of sample.
7 Connections to IV Estimation
Nonadherence has classically been studied in econometrics with IVs. In this work, we focus primarily on the implications of Assumption˜2. Common IV analyses impose additional assumptions such as monotonicity or no defiers; these can impose restrictions in behavior that empirical studies show are sometimes untrue. Nonetheless, in other domains, these assumptions can be more or less plausible, and so, in this section, we discuss connections to other IV models for completeness.
However, a primary point of our paper is that the (potentially covariate-conditional) encouragement effect is the relevant estimand for decision-making purposes. One naive approach to decision-making is the following: Estimate a conditional local average treatment effect with advanced methodology; under monotonicity, compliers drive the encouragement effect, so try to classify who is a complier, and give them the encouragement. An immediate issue here is that complier classification based on the compliance score does not account for the downstream impacts of correct/incorrect classification of compliers and so the classification ought to be weighted by decision-theoretic utility (the clustered local average treatment effect [CLATE]). Making this adjustment just is the conditional encouragement estimand that we focus on in this paper. Further, auditing such a two-stage procedure for fairness introduces complications, as complier classification itself can introduce fair classification concerns. Disaggregating the impacts on takeup/compliance with service enrollment and downstream outcomes can reveal differing sources of inequity and point to different directions for further research.
Fair complier classification auditing under monotonicity

### v5

@@ 5.2 Extension: variance-sensitive constrained policy learning
{assumption}
[Uniform concentration of value and constraint estimators] There exist a scalar sequence $\varepsilon_{V,n}$ and a vector sequence $\varepsilon_{\Delta,n}\in\mathbb{R}_{+}^{K}$ such that, with probability at least $1-\zeta$, and, coordinatewise for each $k\in[K]$,
@@
$\sup_{Q\in\mathcal{P}(\Pi)}\left\{\mathbb{E}[V(Q)]-\mathbb{E}_{n}[v_{DR}(Q)]\right\}\leq\varepsilon_{V,n}(\zeta),\;\;\sup_{Q\in\mathcal{P}(\Pi)}\left\{\mathbb{E}[\Delta_{k}(Q)]-\mathbb{E}_{n}[\delta_{k}(Q)]\right\}\leq\bigl(\varepsilon_{\Delta,n}(\zeta)\bigr)_{k}.$
Under Proposition 5.2 and orthogonal nuisance estimation, $\epsilon_{n}(\zeta)$ follows from standard empirical-process concentration. The optimization approach of Agarwal et al. (2018) follows the no-regret online learning paradigm of Freund and Schapire (1997), an online saddle point optimization alternating optimization over $\lambda$ and best-response updates on $Q$ (hence $\pi$). Further details are in Section 11.2. We only make a small modification to use a refined no-regret (second-order multiplicative weights; Cesa-Bianchi et al. 2007, Steinhardt and Liang 2014) algorithm instead of the original hedge/exponentiated gradient algorithm (Freund and Schapire 1997) for the optimization over $\lambda$. The second-order regret bounds surface dependence on the estimation variance and therefore the two-stage algorithm’s variance reduction.
Two-stage variance-constrained algorithm.
Our two-stage procedure, described formally in Algorithm 1, improves upon generic constrained optimization by introducing implicit variance regularization, so that regret bounds depend more favorably on the maximal variance over small-variance slices near the optimal policy, rather than worst-case constants over all policies (Chernozhukov et al. 2019, Athey and Wager 2021). As a result, such policies can achieve tighter fairness constraint control.
Algorithm 1 Two-stage localized constrained policy optimization
1:Input: policy class $\Pi$, value score $v_{DR}$, vector constraint score $\delta$, target vector $d$, localization radius $\epsilon_{n}$
2:Randomly split the sample into $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$
3:Fit nuisance estimators $\hat{\eta}_{1}$ on $\mathcal{D}_{1}$
4:Compute the first-stage solution $\hat{Q}_{1}\leftarrow\operatorname{REDFAIR}(\mathcal{D}_{1},v_{DR},\delta,d;\hat{\eta}_{1})$
5:Define the nearly binding set $\hat{\mathcal{I}}_{1}\leftarrow\left\{k\in[K]:d_{k}-\mathbb{E}_{n_{1}}[\delta_{k}(\hat{Q}_{1})]\leq\epsilon_{n}\right\}$
6:Compute confidence slacks. For each constraint $\hat{\sigma}_{k}^{2}\leftarrow\widehat{\mathrm{Var}}_{\mathcal{D}_{1}}\!\bigl(\delta_{k}(\hat{Q}_{1})\bigr),$ and set $\hat{d}_{k}\leftarrow d_{k}-z_{\alpha}\hat{\sigma}_{k}\,n_{1}^{-\alpha}$
7:Define the augmented moment and threshold vectors $\tilde{\delta},\tilde{d}$ by
$\tilde{\delta}(Q)\leftarrow\left(\delta(Q)^{\top},\,\bigl(\delta_{\hat{I}_{1}}(Q)-\delta_{\hat{I}_{1}}(\hat{Q}_{1})\bigr)^{\top},\,v_{\mathrm{DR}}(\hat{Q}_{1})-v_{\mathrm{DR}}(Q)\right)^{\top},\qquad\tilde{d}\leftarrow\left(\hat{d}^{\top},\,\epsilon_{n}{\bf 1}_{|\hat{I}_{1}|}^{\top},\,\epsilon_{n}\right)^{\top}.$
8:Fit nuisance estimators $\hat{\eta}_{2}$ on $\mathcal{D}_{2}$. On $\mathcal{D}_{2}$, compute the second-stage solution $\hat{Q}_{2}\leftarrow\operatorname{REDFAIR}(\mathcal{D}_{2},v_{DR},\tilde{\delta},\tilde{d};\hat{\eta}_{2})$}
9:return $\hat{Q}_{2}$
We adapt an out-of-sample regularization scheme developed in Chernozhukov et al. (2019), which recovers variance-sensitive regret bounds. To summarize the procedure at a high level: we split the data into two subsets, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. On the first data split, we estimate the optimal policy distribution $\hat{Q}_{1}$, identify the first-stage binding constraints via the index set $\hat{I}_{1}$, and estimate the objective and constraint variances at $\hat{Q}_{1}$. Next, we augment the constraint matrix to optimize over policy distributions $Q$ in the second stage that are localized within $\epsilon_{n}$-sized neighborhoods for policy value and constraint moment values, relative to $\hat{Q}_{1}$. For example, the new constraint vector slack $\hat{d}_{k}\leftarrow d_{k}-z_{\alpha}\hat{\sigma}_{k}n_{1}^{-\alpha}$ is updated via the binding constraint variance $\widehat{\operatorname{Var}}_{\mathcal{D}_{1}}(\delta_{k}(\hat{Q}_{1})),i\in\hat{I}_{1}$, where $z_{\alpha}$ is the $\alpha$-level z-score to ensure constraints hold with $\alpha$-probability out-of-sample.
$\max_{Q\in\mathcal{P}(\Pi)}$ $\mathbb{E}_{n_{2}}[v_{DR}(Q)]$
s.t. $\mathbb{E}_{n_{2}}[\delta(Q)]\leq\hat{d},\ \mathbb{E}_{n_{2}}\!\left[\delta_{\hat{\mathcal{I}}_{1}}(Q)-\delta_{\hat{\mathcal{I}}_{1}}(\hat{Q}_{1})\right]\leq\epsilon_{n_{2}}\mathbf{1}_{|\hat{\mathcal{I}}_{1}|},\ \mathbb{E}_{n_{2}}\!\left[v_{DR}(\hat{Q}_{1})-v_{DR}(Q)\right]\leq\epsilon_{n_{2}}$
Next, we provide a generalization bound on the out-of-sample performance of the policy returned by the two-stage procedure. Importantly, because of our two-stage procedure, the regret of the policy depends on the worst-case variance of near-optimal policies (rather than all policies). Define the function classes $\mathcal{F}_{\Pi}=\{v_{DR}(\cdot,\pi;\eta)\colon\pi\in\Pi,\eta\in\mathcal{F}_{\eta}\}$, $\mathcal{F}_{j}=\{g_{j}(\cdot,\pi;\eta)\colon\pi\in\Pi,\eta\in\mathcal{F}_{\eta}\}$ and the empirical entropy integral $\kappa(r,\mathcal{F})={\inf_{\alpha\geq 0}\{4\alpha+10\int_{\alpha}^{r}\sqrt{\frac{\mathcal{H}_{2}(\epsilon,\mathcal{F},n)}{n}}{d}\epsilon\}}$, where $H_{2}(\epsilon,\mathcal{F},n)$ is the $L_{2}$ empirical entropy, i.e., log of the $\lVert\cdot\rVert_{2}$ $\epsilon$-covering number. We make a mild assumption of a learnable function class (bounded entropy integral) (Van Der Vaart et al. 1996), which is satisfied by many standard function classes such as linear models, polynomials, kernel regression, and neural networks (Wainwright 2019). {assumption} The function classes $\mathcal{F}_{\Pi},\{\mathcal{F}_{j}\}_{j\in\mathcal{J}}$ satisfy that, for any constant $r,\kappa(r,\mathcal{F})\rightarrow 0$ as $n\rightarrow\infty$. The function classes $\{\mathcal{F}_{j}\}_{j\in\mathcal{J}}$ comprise $L_{j}$-Lipschitz contractions of $\pi.$ We assume that we are using doubly robust/orthogonalized estimation as in proposition 5.4 and, hence, state our results depending on the estimation error of nuisance vector $\eta$. The next theorem summarizes the out-of-sample performance of the two-stage algorithm of Algorithm 1, $\hat{Q}_{2}$.
Theorem 5.6 (Variance-based oracle policy regret)
Let $Q^{\star}\in\arg\max_{Q\in\mathcal{P}(\Pi)}\{V(Q):\Delta(Q)\leq d\}$. Suppose that, with probability at least $1-\delta/2$ over the nuisance function estimation sample, $\max_{\ell\in[L]}\mathbb{E}\!\left[(\hat{\eta}_{\ell}-\eta_{0,\ell})^{2}\right]\leq\chi^{2}_{n,\delta}$ (we have $\chi^{2}_{n,\delta}$ as a bound on nuisance MSE). Denote $v^{0}_{\mathrm{DR}}(O;Q):=v_{\mathrm{DR}}(O;Q,\eta_{0}),\qquad\delta^{0}_{k}(O;Q):=\delta_{k}(O;Q,\eta_{0})$, and define the worst-case variance radius $r:=\sup_{Q\in\mathcal{P}(\Pi)}\sqrt{\mathbb{E}\![(v^{0}_{\mathrm{DR}}(O;Q))^{2}]}.$ For the second-stage sample size $n_{2}:=|D_{2}|$, let $\epsilon_{n_{2}}=\Theta\!(\kappa(r,\mathcal{F}_{\Pi})+r\sqrt{\frac{\log(1/\delta)}{n_{2}}})$. Let $\hat{I}_{1}$ denote the nearly binding set defined in Algorithm 1. Define the localized near-optimal slice
$\mathcal{Q}^{\star}(\epsilon;\hat{I}_{1}):=\left\{Q\in\mathcal{P}(\Pi):V(Q^{\star})-V(Q)\leq\epsilon,\quad\Delta_{k}(Q)\leq d_{k}+\epsilon,\ \forall k\in[K],\quad\Delta_{k}(Q)-\Delta_{k}(Q^{\star})\leq\epsilon,\ \forall k\in\hat{I}_{1}\right\}.$
Let $\tilde{\epsilon}_{n_{2}}=O(\epsilon_{n_{2}}+\chi^{2}_{n,\delta})$. Define the value-variance radius $\bar{\sigma}^{2}_{D_{2}}$ and the constraint-variance:
@@
$\bar{\sigma}^{2}_{D_{2}}:=\sup_{Q,Q^{\prime}\in\mathcal{Q}^{\star}(\tilde{\epsilon}_{n_{2}};\hat{I}_{1})}\operatorname{Var}\!(v^{0}_{\mathrm{DR}}(O;Q)-v^{0}_{\mathrm{DR}}(O;Q^{\prime})),\;\;\bar{\sigma}^{2}_{k,D_{2}}:=\sup_{Q,Q^{\prime}\in\mathcal{Q}^{\star}(\tilde{\epsilon}_{n_{2}};\hat{I}_{1})}\operatorname{Var}\!(\delta^{0}_{k}(O;Q)-\delta^{0}_{k}(O;Q^{\prime}))$
For each $k$, write $\mathcal{F}^{\Delta}_{k}:=\left\{\delta_{k}(\cdot;\pi,\eta):\pi\in\Pi,\ \eta\in\mathcal{F}_{\eta}\right\}.$ Then the policy distribution $\hat{Q}_{2}$ returned by Algorithm 1 satisfies, with probability at least $1-\delta$,
@@
$V(Q^{\star})-V(\hat{Q}_{2})=O\!\left(\kappa(\bar{\sigma}_{D_{2}},\operatorname{conv}(\mathcal{F}_{\Pi}))+\bar{\sigma}_{D_{2}}n_{2}^{-1/2}\sqrt{\log(3/\delta)}+\chi^{2}_{n,\delta}\right),$
and, for each $k\in[K]$,
@@
$\bigl(\Delta_{k}(\hat{Q}_{2})-d_{k}\bigr)_{+}=O\!\left(\kappa(\bar{\sigma}_{k,D_{2}},\operatorname{conv}(\mathcal{F}^{\Delta}_{k}))+\bar{\sigma}_{k,D_{2}}n_{2}^{-1/2}\sqrt{\log(3/\delta)}+\chi^{2}_{n,\delta}\right).$
@@
The specific benefits of the two-stage approach are that 1) the constants are improved from being absolute, structure-agnostic bounds to depending on the variance of low-regret policies, which also reflects the improved variance from the use of doubly robust estimation as in proposition 5.4, and 2) it allows less conservative satisfaction of the fairness constraint out of sample.